OpenStudy (anonymous):

3log_{2}^\left(\begin{matrix}x-1\end{matrix}\right)+log_{4}=4 answer is 3 show me the process please?

OpenStudy (anonymous):

\[3\log_{2}\left( x−1 \right)+\log_{4}=4\]

OpenStudy (anonymous):

no that can't be right. something is missing

OpenStudy (anonymous):

sorry mah bad that second log suppose to be \[3\log_{2}(x-1)+\log_{2}4=4\]

OpenStudy (anonymous):

oop i mean first off \(\log_2(4)=2\) yes?

OpenStudy (anonymous):

so we have \[3\log_2(x-1)+2=4\] and therefore \[3\log_2(x-1)=2\]

OpenStudy (anonymous):

still something is goofy because the answer is not 3

OpenStudy (anonymous):

if you replace \(x\) by 3 you get \[3\log_2(3-1)+2=4\] \[3\log_2(2)+2=4\] \[3+2=4\] and that is not right

OpenStudy (anonymous):

was it \[3\log_2(x-1)+\log_4(4)=4\]??

OpenStudy (anonymous):

then 3 would be the right answer

OpenStudy (anonymous):

\[\log_4(4)=1\] \[3\log_2(x-1)+1=4\] \[3\log_2(x-1)=3\] \[\log_2(x-1)=1\] \[x-1=2\] \[x=3\]

OpenStudy (anonymous):

log_{2}4=5...